# EngGPT2: Sovereign, Efficient and Open Intelligence

### 17 Mar 2026 v1

@@ Abstract
EngGPT2-16B-A3B111[Uncaptioned image]https://huggingface.co/engineering-group/EngGPT2-16B-A3B is the latest iteration of Engineering Group’s Italian LLM and it’s built to be a Sovereign, Efficient and Open model. EngGPT2 is trained on 2.5 trillion tokens - less than Qwen3’s 36T or Llama3’s 15T - and delivers performance on key benchmarks, including MMLU-Pro, GSM8K, IFEval and HumanEval, comparable to dense models in the 8B–16B range, while requiring one-fifth to half of the inference power, and between one-tenth to one-sixth of the training data and consequent needed training power. Designed as a trained-from-scratch Mixture-of-Experts (MoE) architecture, EngGPT2 features 16 billion parameters with 3 billion active per inference, with expert sizes positioned between those used in GPT-OSS and Qwen3. Approximately 25% of its training corpus consists of Italian-language data, to deliver strong capabilities for European and Italian NLP tasks among models of similar scale. This efficiency aims to position EngGPT2 as a key contributor to the growing portfolio of open-weight European models, combining performance and efficiency with full alignment to the EU AI Act. EngGPT2 is also a single model capable of multiple reasoning modes: non-reasoning, reasoning in Italian or English, and turbo-reasoning (a concise, bullet-point style reasoning available in both languages designed for real-time reasoning use cases). EngGPT2 aims to set a new standard for resource-conscious, high-performance LLMs tailored to European and Italian contexts.
@@ 1.1 Context and Scope
Europe is steadily strengthening its capabilities in large language models, even if it still lags behind the major innovation hubs in the United States and China. In recent years, a small but growing set of national initiatives—driven by both industry and academia—has emerged to develop open, region‑focused foundation models. Although still limited in scale, these efforts mark an important starting point for building a European ecosystem committed to responsible, safe, and value‑aligned AI development. As this ecosystem takes shape, generative AI is rapidly becoming a strategic asset for both public administrations and private enterprises. Cultivating strong local expertise is essential: technological sovereignty, sustained innovation, and seamless regulatory compliance all depend on Europe’s ability to design, operate, and govern AI systems within its own principles, standards, and legal frameworks.
Within this context, Engineering Group contributes not only as a technology integrator but as a true solutions builder, combining deep consulting expertise with engineering excellence to guide organizations in defining, designing, and deploying the AI solutions best suited to their specific needs. Our close collaboration with public institutions and industrial partners gives us direct insight into emerging requirements and real‑world operational constraints. This vantage point reinforces the importance of maintaining full control over the entire development pipeline—from initial design to deployment—and of investing in internal capabilities to build, adapt, and continuously refine customized AI systems rather than relying solely on off‑the‑shelf models.
In this spirit, we introduce EngGPT 2, our Large Language Model designed to strengthen Engineering’s technological stack and to contribute to the broader European effort to build a sovereign and open AI ecosystem. Compliance with the EU AI Act is a foundational pillar of this sovereign and open approach. We adopt rigorous transparency practices, including comprehensive technical documentation and the publication of the model on the Hugging Face Hub, to support transparency, accountability, and regulatory compliance. Beyond sovereignty and openness, EngGPT 2 strongly emphasizes efficiency, leveraging advanced architectural choices such as mixture‑of‑experts (MoE), dual‑reasoning flows, and native agentic integrations. These techniques enable significant performance gains—both in speed and in cost of operation—while supporting scalable, adaptable configurations tailored to enterprise and public‑sector environments.
@@ 1.3 Training Compute Assessment and GPAI Classification
The full training pipeline required approximately 250,000 GPU hours, for an estimated total cost of €500,000, assuming an average cost of €2 per GPU hour. Pre-training represented the most computationally intensive stage, accounting for roughly 235,000 A100 GPU hours. This phase ran for approximately 23 days, scaling up to a maximum of 128 nodes (512 GPUs). The observed average Model FLOPs Utilization (MFU) was around 21%, with peak values reaching 31% during the most efficient segments of the run. The long-context training phase required approximately 2,000 GPU hours and was completed in about 20 hours, using up to 128 nodes (512 GPUs). Due to increased sequence lengths and reduced parallel efficiency, the average MFU during this phase was lower, at approximately 14%. The mid-training phase consumed roughly 12,000 GPU hours over a period of 7 days, operating on a reduced configuration of 16 nodes (64 GPUs). The average MFU observed in this stage was approximately 8%, reflecting additional computational overhead introduced by data-dependent processing and masking strategies, which reduce peak hardware utilization in favor of improved training dynamics. Finally, the post-training stages—including fine-tuning and post-training alignment— accounted for approximately 4,000 GPU hours. These activities spanned around 10 days and were executed using variable configurations ranging from 4 to 16 nodes. The MFU during post-training is highly variable across different phases but consistently remains well below 10%. This is expected, as post-training workloads are not compute-bound: they involve autoregressive generation, frequent synchronization, evaluation steps, and small batch sizes, all of which naturally lead to low hardware utilization.
@@ 4.1 Pre-Training Data
Each dataset was used according to its original license and terms of use, and no additional manual annotation was introduced at the pre-training stage. A detailed description of each dataset, including data sources, scale, filtering strategies, and privacy considerations, is provided in Appendix A.
@@
StarCoder Misc [^2] The dataset was filtered to retain only content from repositories identified as open-access / permissively licensed, in accordance with the dataset’s documentation and licensing guidelines (https://huggingface.co/datasets/bigcode/the-stack)) 12% 7% 12%
@@
Table 3: Pretrain datasets

### 30 Mar 2026 v3

@@ Abstract
EngGPT2-16B-A3B111[Uncaptioned image]https://huggingface.co/engineering-group/EngGPT2-16B-A3B is the latest iteration of Engineering Group’s Italian LLM and it’s built to be a Sovereign, Efficient and Open model. EngGPT2 is trained on 2.5 trillion tokens, less than Qwen3’s 36T or Llama3’s 15T, and delivers performance on key benchmarks, including MMLU-Pro, GSM8K, IFEval and HumanEval, comparable to dense models in the 8B–16B range, while requiring one-fifth to half of the inference power, and between one-tenth to one-sixth of the training data and consequent needed training power. Designed as a trained-from-scratch Mixture-of-Experts (MoE) architecture, EngGPT2 features 16 billion parameters with 3 billion active per inference, with expert sizes positioned between those used in GPT-OSS and Qwen3. Approximately 25% of its training corpus consists of Italian-language data, to deliver strong capabilities for European and Italian NLP tasks among models of similar scale. This efficiency aims to position EngGPT2 as a key contributor to the growing portfolio of open-weight European models, combining performance and efficiency with full alignment to the EU AI Act. EngGPT2 is also a single model capable of multiple reasoning modes: non-reasoning, reasoning in Italian or English, and turbo-reasoning (a concise, bullet-point style reasoning available in both languages designed for real-time reasoning use cases). EngGPT2 aims to set a new standard for resource-conscious, high-performance LLMs tailored to European and Italian contexts.
@@ 1.1 Context and Scope
Europe is steadily strengthening its capabilities in large language models, even if it still lags behind the major innovation hubs in the United States and China. In recent years, a small but growing set of national initiatives, driven by both industry and academia, has emerged to develop open, region‑focused foundation models. Although still limited in scale, these efforts mark an important starting point for building a European ecosystem committed to responsible, safe, and value‑aligned AI development. As this ecosystem takes shape, generative AI is rapidly becoming a strategic asset for both public administrations and private enterprises. Cultivating strong local expertise is essential: technological sovereignty, sustained innovation, and seamless regulatory compliance all depend on Europe’s ability to design, operate, and govern AI systems within its own principles, standards, and legal frameworks.
Within this context, Engineering Group contributes not only as a technology integrator but as a true solutions builder, combining deep consulting expertise with engineering excellence to guide organizations in defining, designing, and deploying the AI solutions best suited to their specific needs. Our close collaboration with public institutions and industrial partners gives us direct insight into emerging requirements and real‑world operational constraints. This vantage point reinforces the importance of maintaining full control over the entire development pipeline, from initial design to deployment, and of investing in internal capabilities to build, adapt, and continuously refine customized AI systems rather than relying solely on off‑the‑shelf models.
In this spirit, we introduce EngGPT 2, our Large Language Model designed to strengthen Engineering’s technological stack and to contribute to the broader European effort to build a sovereign and open AI ecosystem. Compliance with the EU AI Act is a foundational pillar of this sovereign and open approach. We adopt rigorous transparency practices, including comprehensive technical documentation and the publication of the model on the Hugging Face Hub, to support transparency, accountability, and regulatory compliance. Beyond sovereignty and openness, EngGPT 2 strongly emphasizes efficiency, leveraging advanced architectural choices such as mixture‑of‑experts (MoE), dual‑reasoning flows, and native agentic integrations. These techniques enable significant performance gains, both in speed and in cost of operation, while supporting scalable, adaptable configurations tailored to enterprise and public‑sector environments.
@@ 1.2 EngGPT 2 Overview
  • Legal Notice: Outlines the copyright terms, usage restrictions, disclaimers, and legal conditions governing the distribution and use of this document.
@@ 1.3 Training Compute Assessment and GPAI Classification
The full training pipeline required approximately 250,000 GPU hours, for an estimated total cost of €500,000, assuming an average cost of €2 per GPU hour. Pre-training represented the most computationally intensive stage, accounting for roughly 235,000 A100 GPU hours. This phase ran for approximately 23 days, scaling up to a maximum of 128 nodes (512 GPUs). The observed average Model FLOPs Utilization (MFU) was around 21%, with peak values reaching 31% during the most efficient segments of the run. The long-context training phase required approximately 2,000 GPU hours and was completed in about 20 hours, using up to 128 nodes (512 GPUs). Due to increased sequence lengths and reduced parallel efficiency, the average MFU during this phase was lower, at approximately 14%. The mid-training phase consumed roughly 12,000 GPU hours over a period of 7 days, operating on a reduced configuration of 16 nodes (64 GPUs). The average MFU observed in this stage was approximately 8%, reflecting additional computational overhead introduced by data-dependent processing and masking strategies, which reduce peak hardware utilization in favor of improved training dynamics. Finally, the post-training stages, including fine-tuning and post-training alignment, accounted for approximately 4,000 GPU hours. These activities spanned around 10 days and were executed using variable configurations ranging from 4 to 16 nodes. The MFU during post-training is highly variable across different phases but consistently remains well below 10%. This is expected, as post-training workloads are not compute-bound: they involve autoregressive generation, frequent synchronization, evaluation steps, and small batch sizes, all of which naturally lead to low hardware utilization.
@@ 4.1 Pre-Training Data
Each dataset was used in accordance with its respective license terms. The training corpus includes a combination of datasets released under open licenses and/or datasets governed by usage-restricted agreements, as well as curated collections of permissively licensed or open-access content. Where datasets included heterogeneous sources (e.g., code or web data collections), additional filtering and selection procedures were applied to retain only content compliant with the applicable licensing terms. A detailed description of each dataset, including data sources, scale, filtering strategies, and privacy considerations, is provided in Appendix A.
@@
StarCoder Misc [^2] The dataset was filtered to retain only content from repositories identified as open-access / permissively licensed, in accordance with the dataset’s documentation and licensing guidelines (https://huggingface.co/datasets/bigcode/the-stack) 12% 7% 12%
@@
Table 3: Pretrain datasets. The listed datasets are subject to different licensing regimes, including open license, usage-restricted agreements and curated collections of permissively licensed content.
@@ 8.1 Model Comparisons
The results obtained with the standard configuration and the optimal configuration are closely aligned, depicting a consistent scenario with only minor variations. Therefore, all subsequent analyses will be conducted based on the optimal configuration.
@@ 9 Conclusion
10 Legal Notice
© Engineering Ingegneria Informatica S.p.A. All rights reserved.
This document is made publicly available for informational and research purposes. The content may be cited, provided that proper attribution is given to Engineering Ingegneria Informatica S.p.A.
No part of this document may be reproduced, modified, or distributed for commercial purposes without prior written authorization from Engineering Ingegneria Informatica S.p.A., unless otherwise explicitly permitted.
This document is provided “as is” and for informational purposes only. Engineering Ingegneria Informatica S.p.A. makes no representations or warranties, express or implied, regarding the accuracy, completeness, or fitness for a particular purpose of the information contained herein.
The content of this document does not constitute any contractual obligation or commitment.